y(y-7)=18
y^2-7y+18=0
from here I couldn't get the middle term
I tried(y-9)(y-2) which is -11; (y-2)(y-9)and got-11 again;(y-18)(y-1)and got-17;(y-1)(y-18)and got -17 again--my book says the answers are -2,9 but I can't get the middle term to break up the equations into the two problems that are needed to get the -2 and 9

Answer by stanbon(75887)

y(y-7)=18
y^2-7y-18=0
---
y^2-9y+2y-18 = 0
y(y-9)+2(y-9) = 0
(y-9)(y+2) = 0
y = 9 or y = -2
==================
Cheers,
Stan H.
=========